OpenStudy (anonymous):

Show that the curve y = 2e^x + 3x + 5x^3 has no tangent line with slope 2. So far, I've set it's derivative to 2... now I'm tempted to take natural logs, but it doesn't seem to lead me anywhere

zepdrix (zepdrix):

\[\Large y\quad=\quad 2e^x+3x+5x^3\]\[\Large y'\quad=\quad 2e^x+3+15x^2\quad=\quad 2\]So this is how far you've gotten, yes? And now we need to show that there are no values of x that will satisfy this. Ok ummm...

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

At least, that's my interpretation too

zepdrix (zepdrix):

If we rearrange things...\[\Large -2e^x\quad=\quad 1+15x^2\] We can think of it like this I suppose, as two separate functions, one on each side. On the right side we have a parabola opening upward from y=1. On the left side we have an exponential function which never goes above y=0. So there is no point of intersection. No solution for x.

zepdrix (zepdrix):

I don't think we can much better than that :( You can't solve for x explicitly since we have the exponential and polynomial x.

OpenStudy (loser66):

I wonder why we have to do more? 3 > 2 , x^2> 0 with any of x, e^x never <0 , the sum of them cannot be 2. It's trivial

OpenStudy (anonymous):

Thanks to both of you

OpenStudy (goformit100):

A warm Welcome to OpenStudy. I can help guide you through this useful site. You can ask your questions to me or you can message me. Please use the chat for off topic questions. Remember to give the person who helped you a medal by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it.